OpenStudy (anonymous):

can somebody please tell me what 4n-n is and how you got the answer?

OpenStudy (hba):

I Think You Can Do It Its A Simple Subraction

OpenStudy (hba):

Its Similar To 4-1

OpenStudy (hba):

\[4girl-1girl=\]

OpenStudy (anonymous):

4n-n is same as 4-1, true that hba :D I doubt it's that simple, though..

OpenStudy (anonymous):

so it would be 3?

OpenStudy (anonymous):

4nin - 1 nins

hartnn (hartnn):

4christinamqualls-1christinamqualls

OpenStudy (anonymous):

your suposed to simplify it it says

OpenStudy (anonymous):

4n-n = n(4-1) = n(3) = 3n

OpenStudy (hba):

You Nailed It Right Anyways You See 4girl-1girl=3girl,Similarly 4n-1n=3n